Price (per hour of tutoring) $25 20 2.50 100 200 300 400 500 600 700 800 900 Quantity (hours of tutoring per week) If the price of tutoring decreases from $15 to $10, producer surplus decreases by, in numerals, $ Your answer

Could someone please show me how to calculate this?

The hint is that the Total Producer Surplus is $3,125...But I do not even know how to calculate that!

I am so lost, please explain!
Answer #1

Producer surplus is the area below equilibrium price and above supply curve.

With P= 15, quantity supplied= 500

Producer surplus= area of triangle formed below the dotted line on 15

Producer surplus= 1/2*(15-2.50)*500= 3125

With P= 10, quantity supplied= 300

New producer surplus= 1/2*(10-2.5)*300= 1125

Change in producer surplus= 3125 - 1125= 2000

producer surplus decreases by 2000